OpenStudy (anonymous):

Cassidy is going to the county fair with her friends but can only stay at the fair for 4 hours. The fair offers two different pricing structures. Cassidy can purchase an all-day pass with unlimited rides for $30.00 or she can pay an admission fee of $6.00 and pay $3.00 for each ride. If Cassidy rides 15 rides, how much does she save by buying the all-day pass?

OpenStudy (anonymous):

$47.00 $27.00 They both cost the same. $21.00

OpenStudy (anonymous):

She would save $21. 3(15)+6=51 51-30=21

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

Ron plans on visiting an amusement park this weekend. Admission to the park costs $11.00, and each attraction costs $2.18 to use. If Ron has saved $50.00 for this weekend, what is the highest number of attractions he can visit at the park?

OpenStudy (anonymous):

18 17 22 4

OpenStudy (anonymous):

He could go on 17 attractions. 50-11=39 39/2.18=17.8899083

OpenStudy (anonymous):

thanks what about this one, The city of Cartesianville is sponsoring an event to collect food for those in need in their community. A local church has already donated 329 pounds of food. The event is expecting 125 attendees. Based on the inequality below, how many pounds of food, f, should each attendee donate in order to collect at least 829 pounds of food? 329 + 125f > 829

OpenStudy (anonymous):

f > 9 f > 4 f > 7 f > 11

OpenStudy (anonymous):

I would say C.

OpenStudy (anonymous):

no that one was A

OpenStudy (anonymous):

A chemist is conducting an experiment which involves a certain chemical reaction. He begins the experiment with 789 milliliters of the chemical, which is consumed by the reaction at a varying rate that is no more than 15 milliliters per minute. Partway through the experiment, the chemist measures the amount of chemical remaining. If there are 669 milliliters of the chemical left, what is the smallest number of minutes that could have passed? 15 minutes 120 minutes 7 minutes 8 minutes

OpenStudy (anonymous):

8 minutes. 15(8)=120 789-120=669

OpenStudy (anonymous):

A division of a company that manufactures widgets has been given an annual operating budget of $68,077.00. The factory's operating costs are determined by two factors - overhead fixed costs of $16,416.39 per year, and a cost of $31.15 to produce each widget. Assuming that the division is constrained by their budget (as in, they cannot exceed their budget), what is the highest number of widgets that the company can produce over the course of the year? 1,658 1,659 400 2,185

OpenStudy (anonymous):

1,658 widgets. 68,077-16,416.39/31.15=1,658

OpenStudy (anonymous):

Marie is saving money for home repairs. To date, she has saved $1,555. She needs at least $2,014 for the repairs. She plans to set aside $51 per week to add to her current savings. If this situation is modeled by the inequality below, how many more weeks, x, does she need to continue saving in order to have enough money for the repairs? $1,555 + $51x > $2,014 x > 21 x > 27 x > 70 x > 9

OpenStudy (anonymous):

thanks for helping also really need help!!!

OpenStudy (anonymous):

You're welcome. :) A? I'm not sure on this one.